Use the linearity property of Laplace transform to find L[5e-2t + t + 2e2t]


1
Expert's answer
2021-09-23T17:15:05-0400

"\mathcal{L}(5e^{-2t} + t + 2e^{2t})\\
=\mathcal{L}(5e^{-2t}) + \mathcal{L}(t) + \mathcal{L}(2e^{2t})\\
=5\mathcal{L}(e^{-2t})+\mathcal{L}(t)+2\mathcal{L}(e^{2t})\\
=\frac{5}{s+2}+\frac{1}{s^2}+\frac{2}{s-2}\\
\frac{5(s^3-2s^2)+s^2-4+2(s^3+2s^2)}{(s^2-4)s^2}\\
=\frac{7s^3-5s^2-4}{(s^2-4)s^2}"